OpenStudy (anonymous):

can anyone factor 6x^2 - 28x + 16?? I am lost because nothing that multiplies into 16 equals 28 when added and 6 isn't a perfect square

OpenStudy (anonymous):

yeah, i can do it

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

first, what two numbers multiple to be 16

OpenStudy (anonymous):

4 4 or 8 2

OpenStudy (anonymous):

more specifically, -4*-4

OpenStudy (anonymous):

right

OpenStudy (anonymous):

but i thought when added together they need to equal -28 that doesnt

OpenStudy (anonymous):

what about 4 and 24

OpenStudy (anonymous):

thats true, but we have to remeber that we have a 6 as the coefficent of the x^2 term, that rule, where they have to add up to the middle term applies mostly when the leading term is x^2 only

OpenStudy (anonymous):

they don't multiply to equal 16

OpenStudy (anonymous):

i am not sure that i know what to do then...

OpenStudy (anonymous):

so what we do is put (6x-?)(x-?)

OpenStudy (anonymous):

-4-24=-28 -4*-24=96

OpenStudy (anonymous):

4 and 4??

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

you serious? its x^2? you had x^4 in the other question...

OpenStudy (anonymous):

now check by foiling out

OpenStudy (anonymous):

(6x-4)(x-4)

OpenStudy (anonymous):

another way you could do it is: 6x2-28x+16 =2(3x2-14x+8) =2(3x2-12x-2x+8) =2{3x(x-4)-2(x-4)} =2(x-4)(3x-2)

OpenStudy (anonymous):

i am sorry junkiejim i didnt catch that... Too many numbers and exponents

OpenStudy (anonymous):

wow, that makes that alot simpler.... here I was trying to figure out how to factor that quartic function....

OpenStudy (anonymous):

its not essential but it would have been somewhat easier if you had first divided the original expression by 2

OpenStudy (anonymous):

i saw that earlier post as well, and knew something off

OpenStudy (anonymous):

I also showed that way

OpenStudy (anonymous):

Thank you very much for the help!!

OpenStudy (anonymous):

you can thank all the people that help you, by clicking the good answer button :)